OpenStudy (anonymous):

1. Solve S = 4v2 for v

OpenStudy (diyadiya):

First Divide 4 on both sides

OpenStudy (diyadiya):

Second ,Take square root on both sides.

OpenStudy (anonymous):

Well first we would divide both sides by 4 and get- S/4=v^2. Now we have to do the square root on both sides sqrt"s/4"=v The s/4 becomes to s/2 and we get Sqrt"s/2"=v

OpenStudy (anonymous):

Answer is \[\sqrt{S/4} = v\]

OpenStudy (anonymous):

Thanks guys<3!!!